# Gamma-ray burst Jet Progenitor: long vs short bursts by angular momentum and mass ratio

### v1

@@ Abstract
In this study, we investigate the homologous collapse dynamics of a slowly rotating stellar nuclear core. We incorporate the virial theorem of hadron collisional relaxations to analyze the conversion of gravitational potential energy into kinetic energy of hadrons. Additionally, we consider the production of photons and electron-positron pairs resulting from hadron collisions. Our analysis reveals a gravo-thermal process wherein gravitational energy transforms into energy associated with photons and pairs. The presence of axial symmetric centrifugal potential reduces the gravitational potential energy gain. The presence of axial symmetric centrifugal potential mitigates the gain in gravitational potential energy. Consequently, we observe the formation of highly energetic and opaque photon-pair jets, whose size, peak spectra, energy and number densities are consistent with observed energetic features and empirical relationships of Gamma-Ray Burst (GRB) progenitors. Furthermore, we find that the angular momentum and mass ratio ($J/M$) of the collapsing core and binary coalescence play crucial roles in determining the jet angle and the differentiation between short and long bursts. Our findings contribute to a deeper understanding of the mechanisms governing GRBs’ progenitor systems.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the most energetic and complex events in the Universe. Within mere seconds, their progenitors release high-energy photons of tremendous energy ranging from $10^{49}$ to $10^{54}$ ergs. Despite significant advancements in our comprehension of GRBs, achieved through extensive research efforts [1, 2, 3, 4, 5, 6, 7], a fundamental question remains unanswered: what are the dominant dynamics of GRB progenitors that ultimately yield the formation of an energetic and optically thick plasma composed of photons and electron-positron pairs, i.e., a fireball or jet ? Massive stellar gravitational potentials are indeed vast reservoirs of energy, and during stellar collapse and coalescence, significant amounts of gravitational energy are released. The critical question is how this immense gravitational energy is converted into photon and pair energies within just a few seconds during GRBs.
In a self-gravitating system, Lynden-Bell and Wood [8, 9] pioneered the study of gravo-thermal catastrophe using the framework of violent hadron collisionless relaxation and the virial theorem. Their work demonstrated the relaxation process in which hadrons gain their mean kinetic heat energy from gravitational potential energy.
On the other hand, the production of photons by hadron collisions has been extensively studied in the context of heavy-ion collisions [10, 11, 12, 13, 14, 15]. Based on the homologous core collapse scenario [16], we show that hadrons gain gravitational energy through collisional relaxation. Hadron-hadron collisions subsequently produce photons, leading to the conversion of gravitational energy into the energies of opaque photon-pair spheres. This process accounts for the origins of GRBs’ energies and their empirical correlations [17].
Furthermore, we investigate the formation of photon-pair jets in scenarios involving rotating core collapse and binary coalescence. We find that the angular momenta and mass ratios of these systems determine the jet angle and the distinct characteristics of long versus short GRBs.
@@
We organize this article as follows. Based on the adiabatic assumption, Sections 2, 3, and 4 present discussions on gravo-thermal catastrophe, the virial theorem, and photon production from hadron relaxations and collisions. In Section 5, we discuss the homologous collapse of a slowly rotating core. Sections 6 and 7 study the properties and correlations of photon-pair spheres and jets. The characteristics of long versus short GRBs, as determined by progenitor angular momenta and mass ratios, are described in Sections 7 and 8.
@@
The stellar core density is characterised by the nuclear saturation density $n_{0}\approx 0.16\,{\rm fm}^{-3}=1.6\times 10^{38}/{\rm cm}^{3}$ or $\rho_{0}=mn_{0}\approx 2.4\times 10^{35}{\rm ergs}/{\rm cm}^{3}$, and $m\approx 1$ GeV is the typical baryon (neutron) mass. The light speed $c=1$, Planck constant $\hbar=1$ and Boltzmann constant $k=1$ are used unless otherwise specified.
@@ 2 Adiabatic approximation in gravitational collapse
Gravitational collapse is a dynamic avalanche involving complex microscopic and macroscopic processes occurring at vastly different scales. It is challenging for both analytical and numerical approaches to analyze all these dynamics without any approximations. To make appropriate approximations, we need to understand the dominant physical processes and their variations across the relevant length and time scales.
@@
Based on this adiabatic approximation, we discuss isothermal configurations during gravitational collapses.
3 Virial theorem and gravo-thermal configuration
We consider a stellar core composed of $N$ baryons of mass $m$ [^1] By the term baryons, we indicate nucleons, nuclei, hadrons, quark-gluon plasma that carry baryon numbers.. In gravitational collapse, baryons gain their mean kinetic energy $F$ from the gravitational energy $U$ via collisionless relaxation. This gravo-thermal catastrophe phenomenon of stellar cores is studied by considering violent relaxation [8] and equipartition theorem [9]. In these studies, the stellar core is an isothermal core of baryon temperature $T\ll m$ and internal “heat” energy $F=\frac{3}{2}TN$. It is an equilibrium or equipartition system, where the Clausius virial theorem applies,
@@ 3 Virial theorem and gravo-thermal configuration
We generalise these discussions to a small fluid element of volume $dV$ inside the stellar core. Given that the baryon-baryon collision timescale at nuclear density is small, equipartition or equilibrium can be locally achieved through the relaxation processes of baryon-baryon collisions. The local virial theorem (3.1) is given by [17]
@@
This relationship shows that via collisional relaxations, baryons gain heat energy (temperature) from the gravitational potential, which is effectively reduced by the centrifugal potential.
4 Photon production from hadron collisions
@@
where $\alpha=1/137$. The QCD coupling $\alpha_{s}\approx 0.5$ indicates that the photon production described by equation (4.2) is dominant over other QED processes producing photons.
@@
The gravo-thermal catastrophe produces not only heat energy for massive baryons, given by $(3/2)T(\rho/m)dV$ (as in equation (3.2)), but also energy for relativistic particles, described by $\rho_{\gamma}dV$ (as in equation (4.2)). For a baryon temperature $T$ on the order of ${\mathcal{O}}(10^{2})$ MeV, the photon density $n_{\gamma}$ (as in equation (4.2)) is so large that relativistic particles become opaque and thermalized by collisions among themselves. The mean temperature $T_{\gamma}$ of thermalized photons can be estimated by equating photon thermal energy density to that given in equation (4.2)
$\frac{\pi^{2}}{15}T_{\gamma}^{4}\approx\rho_{\gamma},\quad\Rightarrow\quad T_{\gamma}\approx 0.21T.$ (4.3)
Moreover, these photons couple to pairs of charged leptons (quarks) and anti-leptons (-quarks). Among these charged particle pairs, we consider pairs of electrons and positrons since their masses are smaller than the photon characteristic energy given in equation (4.3). Therefore, the fourth and back processes involving the production and annihilation of electron-positron pairs, $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, are kinematically possible. The cross-section $\sigma_{\gamma\gamma\leftrightarrow e^{+}e^{-}}$ is of the order of the Thomson cross section $\sigma_{\gamma}$, and the corresponding mean-free length is given by $\sim(\sigma_{\gamma\gamma\leftrightarrow e^{+}e^{-}}n_{\gamma})^{-1}$, which is very small. This indicates that via processes $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, photons and electron-positron pairs participate in thermalisation or thermal equipartition in particle energy and number. As a result, these processes lead to an electrically neutral, deeply opaque, and thermalised plasma (fluid) of photons and electron-positron pairs.
@@
As a consequence, the isothermal core profile is approximately described by baryon temperature $T$ (equation (3.4)) and photon temperature $T_{\gamma}$ (equation (4.3)), as well as the photon production number and energy densities (equations (4.1) and (4.2)). The core will maintain this isothermal profile and evolve homologously.
5 Gravitational collapse of rotating stellar cores
@@
Observations of neutron stars, supernovas, and GRBs suggest that rotation is crucial in the core collapse scenario. However, there is still a lack of realistic models for the final evolutionary state of a massive collapsing star that consistently includes rotation. We proceed along the lines of previous Newtonian and rotating core collapse studies, which construct a set of differentially rotating polytropes in equilibrium as pre-collapse iron core models [20, 21].
@@
We consider a slowly rotating stellar core undergoing a gravitational collapse process, whose rotation energy is much smaller than gravitational energy. In the Newtonian approximation, we describe the process by the continuity equation for the baryon number conservation, Euler’s equation for energy-momentum conservation, Poisson’s equation for gravitational potential $\phi$ are,
@@
5.1 Static equilibrium configuration
To develop a simple analytical model for a rigidly rotating stellar core undergoing gravitational homologous collapse, we start with the established equilibrium conditions and rotation laws. In hydrodynamic equilibrium, the integrability condition of Euler’s equation (5.2) necessitates that the angular momentum $j$ depends only on angular velocity $\Omega$. We adopt the rotation law from Refs. [22, 23, 24],
@@ 5.1 Static equilibrium configuration
$\Omega$$\approx$ $\Omega_{c}/(1+x^{2}/A^{2}),\quad j\approx x^{2}\Omega,$
@@
where $x=R\sin\theta$ and $z=R\cos\theta$ are the cylindrical coordinates, $R$ is the radial coordinate, and $\theta$ is the zenith (polar) angle. The rotating axis is along the $\hat{z}$ direction.
@@
and up to the leading order $\mathcal{O}(x^{2}/A^{2})$,
$j\approx x^{2}\Omega_{c},\quad\phi_{c}\approx\frac{A^{2}x^{2}\Omega_{c}}{2}.$ (5.8)
We set the centrifugal potential to zero at the rotating axis $x=0$, thus ensuring that $\phi_{c}=0$ at $x=0$.
@@
which is uniform across the space.
To generalize this to a model of a rigidly and slowly rotating stellar core undergoing homologous collapse, we need to consider the density, velocity, and potential changes during the collapse. Given the homologous nature of the collapse, the density profile remains similar while scaling with time.

### v2

# Gamma-ray burst collimated fireballs: long vs short bursts by angular momentum and mass ratio

@@ Abstract
In this study, we investigate the gravitational collapses of rotating stellar systems accounting for Gamma-Ray Burst jet progenitors. Based on the virial theorem of hadron collisional relaxations and Newtonian slow-rotating approximation, we analyze the conversion of gravitational binding energy into kinetic energy of hadrons, whose collisions produce photons and electron-positron pairs forming fireballs. Our qualitative analysis implies that rotation effects collimate fireballs with nontrivial angular momenta along the propagating direction, thus possibly making ultra-relativistic jets. Results reveal the possible trends that the fireball becomes more collimated and the jet angle decreases as the total angular momentum and mass ratio $J/M$ of the slow-rotating collapsing core increases. Discussing the extrapolation of these trends to fast-rotating collapsing systems, we speculate that the ratio $J/M$ should be a key quantity for differentiating long bursts (massive core collapses) from short bursts (binary coalescence). We derive the intrinsic correlations of collimated fireball quantities that should be imprinted on a large sample of observed GRB data as empirical correlations.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the most energetic and complex events in the Universe. Within mere seconds, their progenitors release high-energy photons of tremendous isotropic equivalent energy estimated ranging from $10^{49}$ to $10^{54}$ ergs, assuming progenitors originate from opaque spherical fireballs of large particle number and energy densities [1]. The total energy budget should be smaller, considering that GRBs are highly collimated events, with their energy confined within a narrow ultra-relativistic collimated outflow (jet) [2, 3]. The typical jet opening angle is usually a few degrees [4, 5]. For example, the jet opening angle of GRB 170817A is estimated to be around 20 degrees [6, 7]. The jet opening angle of the brightest GRB 221009A is estimated to be exceptionally narrow, approximately 0.6 degrees [8, 9].
Significant advancements in our comprehension of GRBs dynamics and phenomena have been achieved through extensive research efforts [10, 11, 12, 13, 14, 15, 16]. Among others, the ultra-relativistic hydrodynamical outflow of spherical fireballs has been studied well numerically and analytically, e.g., Refs. [10, 15, 17, 18, 19, 20, 21].
However, what are the dominant dynamics of GRB progenitors that ultimately yield the formation of an energetic and optically thick plasma composed of photons and electron-positron pairs, dubbed as a collimated fireball, making an ultra-relativistic jet? There are numerous theoretical investigations in the direction following the Penrose process [22, 23] of extracting rotational energy from particles falling into a spinning black hole (SBH), and Blandford and Znajek’s subsequent work [24, 25] that jet energy could be extracted from the SBH rotational energy and transported out along large-scale magnetic fields surrounding SBH, making a relativistic jet.
In contrast and comparison, we study the possibility of how significant amounts of stellar gravitational binding energy and angular momentum release to collimated fireballs in the gravitational collapse and coalescence of massive and rotating stellar systems. The critical questions are within just a few seconds during gravitational collapses: (i) how this immense gravitational energy is converted (extracted) into photon and pair energies; (ii) why photon and pair energies form opaque and collimated fireballs of huge energy and number densities; (iii) how collimated fireballs acquire nontrivial angular momentum and make ultra-relativistic jets. In Ref. [26], we study how the spherical fireball forms during gravitational homologous collapses of massive stellar cores. Here, considering the collapses of massive and rotating stellar systems, we study the formation of a collimated fireball with nontrivial angular momentum, whose hydrodynamical outflow makes an ultra-relativistic jet.
In a self-gravitating system, Lynden-Bell and Wood [27, 28] pioneered the study of gravo-thermal catastrophe using the framework of violent hadron collisionless relaxation and the virial theorem. Their work demonstrated the relaxation process in which hadrons gain their mean kinetic heat energy from gravitational potential energy. Based on the local virial theorem and the homologous core collapse scenario [29], we show that hadrons gain mean kinetic heat energy from gravitational binding energy through collisional relaxation. On the other hand, the production of photons by hadron collisions has been extensively studied in the context of heavy-ion collisions [30, 31, 32, 33, 34, 35]. These two aspects imply that hadron-hadron collisions subsequently produce photons and pairs, converting their mean kinetic energy to photons’ energy. As a result, the gravitational binding energy converts into the energies of opaque photon-pair plasma (spherical fireball) in gravitational collapses [26]. This article investigates the possibility of forming collimated fireballs and ultra-relativistic jets in gravitational homologous collapses of massive rotating core collapses and binary coalescence.
@@
We organize this article as follows. Based on the adiabatic assumption, Sections 2, 3, and 4 present discussions on gravo-thermal catastrophe, the virial theorem, and photon production from hadron relaxations and collisions. In Section 5, we discuss the gravitational homologous collapse of a slowly rotating core. Sections 6 and 7 study collimated fireball and ultra-relativistic jet properties. Section 8 discusses the intrinsic correlations of collimated fireball properties and how they relate to the empirical correlations from GRB observations. Section 9 discusses the trends of collimated fireball properties as varying progenitor angular momentum and mass ratio, in connection with the characteristics of long versus short GRBs.
@@
The stellar core density is characterised by the nuclear saturation density $n_{0}\approx 0.16\,{\rm fm}^{-3}=1.6\times 10^{38}/{\rm cm}^{3}$ or $\rho_{0}=mn_{0}\approx 2.4\times 10^{35}{\rm ergs}/{\rm cm}^{3}$, and $m\approx 1$ GeV is the typical baryon (neutron) mass. The light speed $c=1$, Planck constant $\hbar=1$, and Boltzmann constant $k=1$ are used unless otherwise specified.
@@ 2 Adiabatic approximation in gravitational collapse
Gravitational collapse is a dynamic avalanche involving complex microscopic and macroscopic processes at vastly different scales. It is challenging for both analytical and numerical approaches to analyze all these dynamics without any approximations. To make appropriate approximations, we need to understand the dominant physical processes and their variations across the relevant length and time scales.
@@
Based on this adiabatic approximation, we discuss local thermal equilibrium configurations during gravitational collapses.
3 Local virial theorem and gravo-thermal configuration
We consider a stellar core composed of $N$ baryons of mass $m$ [^1] By the term baryons, we indicate nucleons, nuclei, hadrons, quark-gluon plasma that carry baryon numbers.. In gravitational collapse, baryons gain their mean kinetic energy $F$ from the gravitational binding energy $U$ via their collisionless relaxation in time-varying gravitational potential. This gravo-thermal catastrophe phenomenon of stellar cores is studied by considering violent relaxation [27] and equipartition theorem [28]. In these studies, authors considered the stellar core to be an isothermal core of baryon temperature $T\ll m$ and internal “heat” energy $F=\frac{3}{2}TN\ll mN$. It is in an equilibrium or equipartition configuration, where the Clausius virial theorem applies,
@@ 3 Local virial theorem and gravo-thermal configuration
We generalise [26] these discussions to a small fluid element of volume $dV$ inside the stellar core by considering a thermal equilibrium locally achieved through the relaxation processes of baryon-baryon collisions at nuclear density. We estimate the baryon collisional relaxation timescale as
$\tau_{\rm relax}=(\sigma_{n}vn)^{-1}\approx 5.21\times 10^{-3}(n_{0}/n)(T/m)^{1/2}T^{-1},$ (3.2)
where $n=\rho/m$ ($v$) is the baryon number density (mean velocity), $\sigma_{n}\approx 1/m_{\pi}^{2}$ is the typical cross-section of baryon-baryon collisions, and averaged collision energy (baryon temperature) $T=(1/2)mv^{2}$. For $T\ll m$, the baryon collisional relaxation time scale $\tau_{\rm relax}$ (3.2) is much smaller than the time scales ($T^{-1}$) of baryon kinetic motion and collisionless relaxation  [^2] In Eq. (3.2), for $T\sim 10^{2}$ MeV and $n\sim n_{0}$, the baryon kinetic motion time scale $T^{-1}\sim 10^{-23}$ seconds, and the baryon collisional relaxation time scale (3.2) $\tau_{\rm relax}\sim 10^{-26}$ seconds, i.e., $\tau_{\rm relax}\ll T^{-1}$.. Therefore, a local thermal equilibrium can establish, and the Clausius virial theorem applies. Analogously with global virial theorem (3.1), the local virial theorem is given by
@@
This relationship shows that through collisional relaxations, baryons gain heat energy (temperature) from the gravitational potential, which is effectively reduced by the centrifugal potential.
4 Photon and $e^{+}e^{-}$ pair production from hadron collisions
@@
where $\alpha=1/137$. The QCD coupling $\alpha_{s}\approx 0.5$ indicates that the photon production described by equation (4.2) dominates other QED processes producing photons.
@@
The gravo-thermal catastrophe produces not only heat energy for massive baryons, given by $(3/2)T(\rho/m)dV$ (as in equation (3.3)), but also thermal energy for relativistic particles, described by $\rho_{\gamma}dV$ (as in equation (4.2)). For a baryon temperature $T$ on the order of ${\mathcal{O}}(10^{2})$ MeV, as will be shown later, the photon number density $n_{\gamma}$ (4.1) and energy density $\rho_{\gamma}$ (4.2) are very large. It implies that photons become opaque and thermalized by interacting with pairs of electrons and positrons. The local mean temperature $T_{\gamma}$ of thermalized photons can be estimated by equating photon thermal energy density $T^{4}_{\gamma}$ to the energy density $\rho_{\gamma}$ given in equation (4.2)
$\frac{\pi^{2}}{15}T_{\gamma}^{4}\approx\rho_{\gamma},\quad\Rightarrow\quad T_{\gamma}\approx 0.21T,$ (4.3)
because the local thermal equilibrium is approximately established. This result shows that the local mean photon temperature $T_{\gamma}$ is almost the same order as the local mean baryon temperature $T$.
The fourth and back processes involving the production and annihilation of electron-positron pairs, $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, are kinematically possible because the photon characteristic energy scale $T_{\gamma}$ (4.4) is larger than electron and position masses $2m_{e}$. The cross-section $\sigma_{\gamma\gamma\leftrightarrow e^{+}e^{-}}$ is of the order of the Thomson cross-section
$\sigma_{\gamma}\approx\sigma_{T}(3/8)(m_{e}/T_{\gamma})\ln(2T_{\gamma}/m_{e})$
and $\sigma_{T}=(8\pi/3)\alpha^{2}/m^{2}$. The corresponding photon mean-free length is given by
$\lambda_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}\approx\frac{3.23\times 10^{6}}{[\ln(2T_{\gamma}/m_{e})]}\left(\frac{m_{e}T_{\gamma}}{T^{2}}\right)\frac{1}{T}\sim{\mathcal{O}}(10^{-10})~{}{\rm cm},$ (4.4)
which is much smaller than a macroscopic size, namely optical depth $\tau_{\gamma}$ is very large.
This indicates that via processes $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, photons and electron-positron pairs participate in thermalisation or thermal equipartition in particle energy and number. As a result, these processes lead to an electrically neutral, deeply opaque, and thermalised plasma (fluid) of photons and electron-positron pairs, i.e., a “collimated fireball”. More detailed discussions can be found in Ref. [26].
@@
As a consequence, the isothermal core profile is approximately described by baryon temperature $T$ (equation (3.5)) and photon temperature $T_{\gamma}$ (equation (4.3)), as well as the photon production number and energy densities (equations (4.1) and (4.2)). The core will maintain these isothermal profiles and evolve homologously.
5 Gravitational collapse of slowly rotating stellar cores
@@
Observations of neutron stars, supernovas, and GRBs suggest that rotation is crucial in the core collapse scenario. However, realistic models for the final evolutionary state of a massive collapsing star that consistently includes rotation are still lacking. We proceed along the lines of previous Newtonian and rotating core collapse studies [38, 39], which construct a set of differentially rotating polytropes in equilibrium as pre-collapse iron core models.
5.1 Euler and Poisson equations in Newtonian approximation
We consider a slowly rotating stellar core undergoing a gravitational collapse process, whose rotation energy is much smaller than gravitational binding energy. Following the semi-analytical approach [29, 26], we will study the homologous gravitational collapse of slowly rotating and massive spherical cores by using the gravitational and centrifugal potentials in Newtonian approximation. We aim to explain how hadron collisions in gravitational potential balanced by centrifugal potential convert gravitational binding energy to the thermal energy of photons and pairs, creating a collimated fireball during gravitational collapses. Microscopic physical processes, e.g., photon production by hadron collisions, at a small spacetime scale are not affected much by using the general relativity (GR) or Newtonian framework at a macroscopic spacetime scale. However, it affects the spacetime distributions of collimated fireball properties, e.g., energy density, particularly approaching horizons. In this study, we attempt to obtain the approximate results and qualitative estimations of the total collimated fireball thermal energy, size and other quantities averaged over the entire collimated fireball volume and formation time at the end of collapses. We also consider that such approximate results should indicate a lower bound of counterpart energetic quantities obtained in the GR framework since the GR gravitational binding energy (potential) is larger than the Newtonian one in the gravitational collapse process from infinity to horizon $2M$. The numerical approach based on the GR framework is necessary for quantitative studies of collimated fireball formation from gravitational collapses of massive rotating stellar and binary systems. It is fundamental for the fireball’s inner part dynamics near the horizon. Nevertheless, due to the complex physical issues of GRB progenitors, we first figure out what macroscopic and microscopic processes are dominant and crucial for producing and collimating fireballs. Therefore, it deserves preliminary analysis using the analytical approach based on the Newtonian approximation in the slow-rotating limit.
@@ 5.1 Euler and Poisson equations in Newtonian approximation
In the Newtonian approximation, we describe the process by the continuity equation for the baryon number conservation, Euler’s equation for energy-momentum conservation, Poisson’s equation for gravitational potential $\phi$ are,
@@
5.2 Static equilibrium configuration and rotation law
We start with the established equilibrium conditions and rotation laws to develop a simple analytical model for a rigidly rotating stellar core undergoing gravitational homologous collapse. In hydrodynamic equilibrium, the integrability condition of Euler’s equation (5.2) necessitates that the angular momentum $j$ depends only on angular velocity $\Omega$. We adopt the rotation law from Refs. [40, 41, 42],
@@ 5.2 Static equilibrium configuration and rotation law
$\Omega$$\approx$ $\Omega_{c}/(1+x^{2}/A^{2}),$ (5.6)
where $x=R\sin\theta$ and $z=R\cos\theta$ are the cylindrical coordinates, $R$ is the radial coordinate, and $\theta$ is the zenith (polar) angle. The rotating axis is along the $\hat{z}$ direction. The system is axially symmetric and we omit the coordinate $y$ in formulae. The centrifugal potential is given by
@@
and $d\phi_{c}=jd\Omega$.
@@
and up to the leading order $\mathcal{O}(x^{2}/A^{2})$, the rotation law becomes
$j\approx x^{2}\Omega_{c},\quad\phi_{c}\approx\frac{1}{2}x^{2}\Omega^{2}_{c}\approx\frac{1}{2}j\Omega_{c}.$ (5.9)
Here, we drop the constant in Eq. (5.7) by ensuring the centrifugal potential to zero $\phi_{c}=0$ at the rotating axis $x=0$. We note that $hdm$ and $\phi dm$ are elementary heat and gravitational binding energies, $jdm$ and $\phi_{c}dm$ are elementary angular momentum and rotation (kinetic) energy for a baryon fluid element of mass $dm=\rho dV$ inside a rotating core.
@@
which is uniform across the space. The constant on the right-handed side relates to the material binding energy. To generalize this to a model of a rigidly and slowly rotating stellar core undergoing homologous collapse, we need to consider the density, velocity, and potential changes during the collapse.